### v1

@@ 2 Results
Figure 5: Comparisons between human participators and LMMs for the scientific figure understanding task. (a) Boxplots of average scores across all datasets by different participators. (b) Boxplots of participators’ performances by different problem categories. (c) Boxplots of accuracy by confidence levels provided by different human participators. We have three options for humans to choose. (d) The first case study is based on the original question, LLM-generated answer, and Human-generated answer. (e) The second case study is based on the original question, LLM-generated answer, and Human-generated answer.
@@
Interpretation of the Human-Eval results of figure understanding. In this part, we summarize the performances of human participators for the figure understanding task and compare the differences and similarities between human-generated answers and LMM-generated answers. To examine the upper bound of LMM performance, we included the 5 major voting with o1, o1 (5 mv), as a new baseline and the best-of-5 with o1, o1 (bo5), as the oracle for o1. We selected eight human experts, numbered from H1--H8, to conduct experiments. According to Figure 5 (a), the performer with the best average accuracy in answering figure-related questions is still human, while LMMs with test-time computation with three different variations can only surpass two human participators. Moreover, major voting improves the accuracy of the LMM and reduces the difference in performance compared with the best human participators. This conclusion is supported by the p-value difference of the two-sided Wilcoxon rank-sum test. However, major voting consumes more tokens and therefore leads to higher costs. Therefore, we also explored how to utilize rationale engineering to improve the answer accuracy of LMM in the next section. In Figure 5 (b), we visualize the differences between the performances of LMMs and humans by categories. We found that o1 (5 mv) ranks higher than o1 in figure summarization and counting ability apparently, the LLMs might have more randomness in generating answers related to these tasks. Moreover, in all of the categories, the best performer is human, but the human participators also have in-group variation across different categories. However, according to Extended Data Figure 3 (b), LMMs still have advantages in reducing both the time cost and monetary cost for this task. Here the monetary cost of LMMs is computed based on the token price of API calling [61]. Therefore, if we want to utilize the responses of human participators to improve the performance of the LMM, we need to consider the reference population in detail and select human participators with higher correctness rates to generate the data.
@@
When designing the questionnaire, we also asked human participators to provide confidence evaluation for their outputs, stratified by I don’t know (in short, idk, which can be interpreted as the lowest certainty or random guess), low certainty (low), and high certainty (high). According to Figure 5 (d), the accuracy from the high group is significantly larger than the accuracy from other groups, validated by the p-value from the Mann-Whitney U test. Therefore, we can conclude that the confidence given by humans is more accurate than LLM-generated confidence, and thus the human-generated confidence level might be helpful for better calibrating the LLM outputs.
Here we also provide two examples for the error analysis of generated answers. We define the error types by following related work[62, 63]. In Figure 5 (d), we present a question where all of the LLMs and most of the humans made the wrong decision. This question required participators to describe the tendency between CPS value and relative abundance of different cell types across two regions. Here we found that o1 tended to group the two different regions together and presented the shared patterns, while the correct answers generated by the human participators separated the analyses for both regions and cell types. As a result, LLMs might overlook specific details, and their ability to understand the details of a picture and make inferences needs to be strengthened. The main reason that human participators made wrong decisions in this question is that they did not separate the consideration by different cell types, which also reflects the hardness of extracting effective information in the scientific figure. We also present one example where all of the LMMs made the wrong decision but nearly all of the human participators could give the correct answer, shown in Figure 5 (e). This question required participators to identify the method with the largest variation across different metrics. Here we found that LMMs selected the wrong option even after checking the width of box-plot distributions, which means that the reasoning path is correct. Human participators tended to directly give the answer, but the reasoning path should be similar to the LMMs’ options as we need to recall the definition of the width of boxplots to address this question. All of the error types can be found in Supplementary File 1.
Moreover, we conducted the analysis of error sources for both LMMs and human participators, which can be classified into three different classes: 1. reasoning process error; 2. extract information error; and 3. generation error. According to Extended Data Figures 6 (a) and (b), the error with type 2 has the largest proportion for both LMMs and human participators, and the error with type 1 sometimes occurred. Human participators do not have type 3 errors, which demonstrates a stronger ability to follow instructions. Moreover, LMMs with smaller scales might also have generation errors, reflected in missing outputs or out-of-context generation. We also explored the understanding gap between o1 and the best human participator by compressing their outputs from text space into embeddings space, and the cosine similarity between the embeddings using all the selected questions from different figures is shown in Extended Data Figure 4 (b). This figure suggests that there still exist some differences between the output produced by a human and that by an LMM, which inspired us to explore the use of reasoning techniques to enhance human understanding of scientific figures with the help of o1-type models. Therefore, we observe the alignment and differences between LMMs and humans in interpreting the difficult questions, which are valuable resources to guide future work on improving the figure understanding ability of LMMs with human feedback or outputs, or conversely.
@@
Figure 6: Results of verification-correction LMMs as collaborators for helping human researchers. (a) Landscape of proposed self-correction pipeline by ARIEL. The solid arrows in the panel indicate the steps that must be executed, and the dashed arrows indicate the steps that are optional. (b) Comparisons of performances among human participators and human participators with ARIEL (self-corrected LMMs) under two different modes. (c) Comparisons of performances among human participators, human assistants, and ARIEL. We averaged the scores of six human assistants to compare. (d) Sankey plots for demonstrating the improvement of using ARIEL as a research assistant for human experts H2 and H7. We count the number of questions based on the conditions of with/without assistance and correctness by setting 0.5 as the accuracy threshold. (e) The performances of different options by problem categories. (f) One case study with wrong solutions coming from all options. (g) One case study with wrong solutions from humans, which is later corrected by LMMs.
@@
Verification and Correction LMMs help humans understand scientific figures. In this section, we demonstrate that using LMMs to conduct verification and correction can correct humans’ incorrect understanding of selected figures with both qualitative and quantitative performance. The pipeline is summarized in Figure 6 (a), and the rest of the figures present our results. Figure 6 (b) shows the accuracy of three participators (labeled H1, H2, and H7) under three conditions: the solution written by themselves without LLM assistance (pink bar), the solution is verified and corrected by an LLM assistant using our verification-correct framework (yellow bar for the default mode and green bar for the uncertain mode). Accuracy (ranging from about 40.0 to 100.0) is plotted on the vertical axis, while each participator’s results are grouped along the horizontal axis. Comparing the Raw (pink) boxes, H1’s baseline accuracy is clearly higher than that of H2 and H7. When the LLM assistant (yellow) or LLM assistant uncertain (green) is applied, the corrected results of H2 and H7 (they originally performed at a lower accuracy) show measurable improvements. Statistically significant p‐values shown in the same figure indicate that these gains are meaningful. Overall, our results illustrate that human participators with lower initial analytical performance benefited notably from LMM assistance, improving their answer accuracy by a large margin. This observation is further supported by comparing the performances between ARIEL and human assistants for correcting the answers from participators, shown in Figure 6 (c) and Extended Data Figure 7, which demonstrates that ARIEL has a generally better performance compared with the average level of human experts. Meanwhile, we show the improvement of using ARIEL as a corrector versus the raw answers via Sankey plots for H2 and H7. Based on Figure 6 (d), ARIEL can improve the accuracy of answers for over 65% questions with low initial accuracy, further supporting the strong capacity of ARIEL as a research assistant.
@@
Figure 6 (e) shows a breakdown analysis over different categories of biomedical questions, including figure summarization, insight discovery, color identification, model design, counting ability, and rank identification. Applying an LLM assistant over H2 or H7 consistently outperformed H2 and H7 for all categories, indicating the effectiveness of the LLM assistant for verifying and correcting H2 and H7 solutions. Applying an LLM assistant in the uncertainty mode where “I am not sure about the answer” is added after the human answer to deliberately encourage LLM to correct the answer outperforms the original H2 and H7 performance in most categories except for rank identification. Our LMM output inspection shows that LLM generates relevant context, but it struggles to consistently produce a final ranked answer for the rank identification task when the uncertainty mode is enabled. Among all categories, the accuracy of answers for questions belonging to insight discovery shows the largest variance across both human participator answers and LMM assistant answers, which is consistent with our results discussed in the previous sections. This indicates that scientific discovery is constrained by the capabilities of human experts and remains challenging even with the assistance of an LMM.
We present case studies in Figures 6 (f) and (g) to further illustrate our results. In Figure 6 (f), while the human participator was unable to identify the desired panel from the figure, an LMM assistant also provided a wrong answer, but it made a more effortful attempt than the human participator by providing the rationales. In Figure 6 (g), while the human participator was confused about the figure due to its complexity, the LLM assistant was able to retrieve the relevant knowledge and identify, extract, and explain the patterns of the region in “E”s from the figure to correctly answer the question. Since we allowed human participators to search for external resources to answer this question, our finding implies that LMM has a stronger ability to integrate knowledge about genomic annotation to answer this question. Furthermore, the LMM can also correct the incorrect answer made by the human participator, implying that our proposed pipeline can fix the incorrect answer caused by the knowledge gap effectively. In summary, our LMM verification and correction framework successfully enhances the performance of less analytically skilled human participators, bringing them to a level comparable to the top experts.
@@
Figure 7: Landscape and results of generating hypothesis from multimodal inputs. (a) Our proposed framework of generating scientific hypothesis with two different modes. (b) Evaluation results of different generated outputs with four metrics and their averaged score (Avg). (c) Examples of generated hypotheses with the same image inputs but from different modes.
@@
Exploring the hypothesis-generation capacity of ARIEL framework with text-image paired inputs. Finally, we investigate the possibility of leveraging ARIEL with the understanding of multimodal information to generate scientific hypotheses to inspire novel ideas in biomedical research. Here we use scientific images from the previous question but with new query prompts focusing on hypothesis generation. Moreover, we consider two different modes to generate the hypothesis, including the single-LLM mode with o1 as the base model and the Multi-LMM agent mode (HypoGenAgent) with o1 as the base model. HypoGenAgent is a system with four different LMMs with different roles, modified from [29]. The overview of our proposed method is summarized in Figure 7 (a). To evaluate the quality of the generated hypothesis, we consider different aspects modified from [64] with a comprehensive cover for evaluating scientific insights:
@@
Each of the assessment criteria here can be quantified as a score between 0 and 100, where 0 means not at all compliant and 100 means fully compliant. To reduce the bias introduced in the evaluation process, we utilize Claude 3.7 Sonnet as a judge [65] and score each output from different modes, shown in Figure 7 (b). According to this figure, although HypoGenAgent (o1) has a higher score compared with o1 on average, we do not observe statistically significant differences in the hypothesis generated by two different modes (p-value$>0.05$ with two-sided Wilcoxon rank-sum test). To better understand the properties of the hypothesis generated by these modes, we select representative examples of generated hypotheses for in-depth analysis. Based on Figure 7 (c), we found that the hypothesis from o1 generally has a longer length and more figure-oriented information, while HypoGenAgent (o1) tends to generate hypotheses with shorter lengths and more general plans. Therefore, there exists a completeness-conciseness balance when using ARIEL as a hypothesis generator, which may help researchers to design ideas that fit well with their requirements.
@@
  • With the help of verification and correction strategy, we can leverage the reasoning ability of advanced LMMs to help underperforming human participators by correcting their wrong answers, and thus improve their performances. Therefore, our experiment shows the potential of using LMMs as assistants for understanding complicated scientific figures.
@@ 3 Discussion
Furthermore, our results based on difficulty analysis imply that LLMs even have consistency with humans in their understanding of difficult problems. For example, LLMs also usually summarized abstract with low scores on the articles that most human participators marked as most difficult. Meanwhile, the answers for questions with the insight discovery category in the scientific figure understanding task also had larger variation compared with other categories, which was observed in both outputs from humans and outputs from models. These findings provide a new perspective on generation errors in foundation models. Finally, we demonstrated that ARIEL can also work as a hypothesis generator with multimodal inputs for accelerating scientific discovery, which aligns well with the perspectives of having agentic AI for accelerating scientific research [66].
@@ 4 Methods
Problem Statement. In this manuscript, we mainly focus on evaluating the outputs of LLMs understanding biomedical manuscripts and comparing the human- and LLMs-generated abstracts or answers. Here we define the participator $\mathcal{M}$ in the experiment as $O=\mathcal{M}(P,T,I)$, where $O$ represents outputs, $P$ represents prompts, $T$ represents text inputs, and $I$ represents image inputs. To evaluate the quality of LLMs-generated abstracts, $I$ is placed as empty, and we treat the original abstracts as ground truth $O^{T}_{G}$ and compare their differences by both quantitative approaches (denoted as evaluator $E(O_{G},O)$) and qualitative approaches. We also invite human experts to summarize the same testing documents and compare the differences from outputs. To evaluate the quality of LMMs-generated answers for figure understanding, $I$ is paired with question $P$, and $T$ is empty. Here we compare their outputs with human-annotated ground truth $O^{I}_{G}$ and invite human experts to answer the same questions and evaluate both human- and LMMs-generated outputs.
@@
Regarding the first part of the evaluation dataset, we download the open-access research manuscripts from PubMed [31] with a knowledge cut-off date later than 2024 January 1st and earlier than 2024 June 1st. The total dataset contains $n=2000$ papers, covering different topics. We further split the papers into training and testing datasets with a ratio of 0.7/0.3. The training dataset was used to fine-tune specific models, while the testing dataset was used to evaluate both open-source and closed-source models. For one paper, we took the abstract section out as ground truth, and deleted this section in the original paper as input texts. Our instruction was used to query the model to give us the summarization of targeted research manuscripts.
Regarding the second part of the evaluation dataset, we collected ten scientific figures from published manuscripts in top-tier journals focusing on genomics and proteomics research [23, 36, 37]. We designed several questions to evaluate the performances of LMMs with different aspects, including figure summarization, color identification, model design, rank identification, recommendation, and insight discovery. We adjusted the prompt to provide necessary background information with questions as instruction and input texts, and combined the prompt with the paired figure as inputs of different LMMs. We generated the ground truth information of each question by referring to the original description of these figures in the manuscript.
@@
For the long document summarization task, we consider seven different metrics for evaluation, including BLEU, ROUGE-1, ROUGE-2, ROUGE-L, BERT score (BERT), F1-Radgraph, and MEDCON, for evaluation [48, 49, 50, 51, 47]. Details of the metrics are introduced below:
@@
5. MEDCON: MEDCON restricts the concepts and entities in QuickUMLS [67] semantic groups (Anatomy, Chemicals & Drugs, Device, Disorders, Genes & Molecular Sequences, Phenomena and Physiology) and we extract such sets from generated texts and ground truth texts. We then compute the F1 score based on these two sets. This score ranges from 0 to 1 and is later scaled to (0,100). A higher score means better model performance.
Based on the evaluation results recorded in [12], BERT and MEDCON scores have a high correlation with the correctness of outputs, while BLEU and ROUGE scores have a high correlation with the completeness of outputs. We compute the final score of each method by averaging the metric scores across all the testing manuscripts. The scores are scaled all into 0-100 and higher scores mean better model performances.
For the scientific figure understanding task, by following the design of [25, 23], we only consider the accuracy by comparing model outputs and ground truth information. Since we also need to consider the inference information as well as format information of model outputs, and thus our previous settings for long document summarization evaluation cannot be directly applied. We utilize Spacy [68] as a tool to split model outputs into different sentence blocks with similar methods and evaluate the correctness of each block. The accuracy is computed based on the number of correct blocks and the total number of blocks, for each question. For each figure, we compute the average score across different questions to report the performance of specific models in the given figure.
@@
LMM verification and correction to improve figure understanding capacity of humans. Scaling LMM test-time compute has been shown to be effective in improving the reasoning performance of LMMs on math and logical reasoning[69, 59, 70]. We adopt the self-verification and self-correction component of SETS [69], a test-time scaling framework for iterative verification and correction and propose a verify-correct framework using an LMM to conduct verification and correction over human-written solutions for biological figure reasoning problems. In contrast with SETS which adopt self-verification and self-correction over LLM outputs, we conduct verification and correction over human outputs as well as iterative self-verification and self-correction over LMM-generated outputs.
@@
We use one of the most advanced general-purpose Multimodal reasoning models, OpenAI o1 [58] for evaluating LLM verification and correction over improving figure understanding capacity of humans. The temperature is set to 1 to promote the model to generate diverse and creative reasoning steps. The maximum iterations of verification and correction are set to 10 by following the general setting in this method [69] (In our task, observed minimal iteration is 6, and maximal iteration is 10). The workflow is illustrated in Algorithm 1.
@@
1:Input: System prompt $T_{S}$, question $Q_{M}$, scientific image $I_{S}$, human outputs $O_{H}$, verify prompt $T_{V}$, correct prompt $T_{C}$, number of iteration $N$.
2:Helper Models: An LMM $\mathcal{M}(\cdot,\cdot,\cdot)$, concatenation function |$\cdot$|$\cdot$ | | ⋅.
@@
We propose two different modes for this process, including default mode and uncertain mode. The uncertain mode means we add “I am not sure about the answer.” at the end of each human output to let the LMM agent be aware of the uncertainty level of such answers.

### v2

@@ 2 Results
Figure 5: Comparisons between human participators and LMMs for the scientific figure understanding task. We report average and standard error in all bar plots. (a) Bar plot of average scores across all datasets by different participators. (b) Bar plots of accuracy and standard errors from participators’ performances by different problem categories. (c) Bar plot of accuracy by confidence levels provided by different human participators. We have three options for humans to choose. (d) The first case study is based on the original question, LLM-generated answer, and Human-generated answer. (e) The second case study is based on the original question, LLM-generated answer, and Human-generated answer.
@@
Interpretation of the Human-Eval results of figure understanding. In this section, we compare the performance of human participants and LMMs on the figure understanding task, highlighting both similarities and differences in their outputs. To approximate the upper bound of LMM performance, we introduced two additional baselines for o1: the five-majority-vote variant (o1 (5 mv)) and the best-of-five variant (o1 (bo5)), the latter option serving as an oracle. We also had eight human experts (H1–H8) in the evaluation. According to Figure 5 (a), the performer with the best average accuracy in answering figure-related questions is still human, while LMMs with test-time computation with three different variations can only surpass two human participators. Moreover, major voting improves the accuracy of the LMM and reduces the difference in performance compared with the best human participators. This conclusion is supported by the p-value difference of the two-sided Wilcoxon rank-sum test. However, major voting consumes more tokens and therefore leads to higher costs. Therefore, we also explored how to utilize rationale engineering to improve the answer accuracy of LMM in the next section. In Figure 5 (b), we visualize the differences between the performances of LMMs and humans by categories. We found that o1 (5 mv) ranks higher than o1 in figure summarization and counting ability apparently, the LLMs might have more randomness in generating answers related to these tasks. Moreover, in all of the categories, the best performer is human, but the human participators also have in-group variation across different categories. However, according to Extended Data Figure 3 (b), LMMs still have advantages in reducing both the time cost and monetary cost for this task. Here the monetary cost of LMMs is computed based on the token price of API calling (62). These findings indicate that integrating human feedback remains valuable for improving LMM performance. However, to maximize benefits, careful selection of reference participants is necessary, prioritizing individuals with higher accuracy to ensure the reliability of training or evaluation data.
@@
When designing the questionnaire, we also asked human participators to provide confidence evaluation for their outputs, stratified by I don’t know (in short, idk, which can be interpreted as the lowest certainty or random guess), low certainty (low), and high certainty (high). According to Figure 5 (d), responses in the high-confidence group achieved significantly higher accuracy than those in the other groups, with statistical significance confirmed by the Mann–Whitney U test. These results suggest that human self-assessed confidence is more reliable than LLM-generated confidence and may serve as a useful signal for calibrating LLM outputs.
We further conducted error analysis of generated answers, with error types defined following prior studies (63, 64). Figure 5 (d) illustrates a case in which all LLMs and most human participants produced incorrect responses. The task required describing the relationship between CPS values and the relative abundance of cell types across two regions. In this instance, o1 grouped the two regions together and emphasized shared patterns, whereas the correct human responses separated analyses by both region and cell type. This outcome suggests that LLMs often overlook fine-grained details and require improved capacity for figure-specific inference. Human errors in this example primarily stemmed from not stratifying the analysis by cell type, highlighting the intrinsic difficulty of extracting precise information from scientific figures. A contrasting example is shown in Figure 5 (e), where all LLMs failed but nearly all human participants answered correctly. The question asked participants to identify the method with the greatest variation across metrics. While LLMs incorrectly selected another option, their reasoning process appeared consistent with human logic, as both relied on interpreting boxplot widths. The discrepancy suggests that although LLMs followed the correct reasoning path, they failed in execution, whereas humans directly identified the correct method. All identified error categories are summarized in Supplementary File 1.
We further analyzed the error sources for both LMMs and human participants, categorizing them into three classes: (1) reasoning process errors, (2) information extraction errors, and (3) generation errors. As shown in Extended Data Figures 7 (a) and (b), type 2 errors accounted for the largest proportion in both groups, while type 1 errors occurred less frequently. Notably, human participants did not exhibit type 3 errors, underscoring their stronger ability to follow instructions. By contrast, smaller-scale LMMs occasionally produced generation errors, such as missing outputs or out-of-context responses. To further assess the alignment between humans and models, we compared the outputs of o1 with those of the best-performing human participant by projecting them into an embedding space and calculating cosine similarity (Extended Data Figure 5 (b)). The results indicate residual differences between human- and LMM-generated outputs, suggesting an incomplete overlap in their interpretive processes. This observation motivates the exploration of reasoning-enhancement strategies, leveraging o1-type models to support human interpretation of scientific figures. Overall, these findings highlight both the alignment and divergence between LMMs and humans in addressing complex figure-based questions, offering valuable insights for advancing LMM figure comprehension through human feedback and, conversely, for augmenting human analysis with model-assisted reasoning.
@@
Figure 6: Results of verification-correction LMMs as collaborators for helping human researchers. We report the average and standard error in all bar plots. (a) Landscape of proposed self-correction pipeline by ARIEL. The solid arrows in the panel indicate the steps that must be executed, and the dashed arrows indicate the steps that are optional. (b) Bar plots of comparisons of performances among human participators and human participators with ARIEL (self-corrected LMMs) under two different modes. (c) Bar plots of comparisons of performances among human participators, human assistants, and ARIEL. We averaged the scores of six human assistants to compare. (d) Sankey plots for demonstrating the improvement of using ARIEL as a research assistant for human experts H2 and H7. We count the number of questions based on the conditions of with/without assistance and correctness by setting 0.5 as the accuracy threshold. (e) Bar plots of accuracy and standard errors from different options by problem categories. (f) One case study with wrong solutions coming from all options. (g) One case study with wrong solutions from humans, which is later corrected by LMMs.
@@
Verification and Correction LMMs help humans understand scientific figures. In this section, we demonstrate that LMMs can be effectively employed for verification and correction, thereby improving human interpretation of scientific figures in both qualitative and quantitative terms. The overall pipeline is summarized in Figure 6 (a), with subsequent panels presenting the results. Figure 6 (b) shows the accuracy of three participators (labeled H1, H2, and H7) under three conditions: the solution written by themselves without LLM assistance (pink bar), the solution is verified and corrected by an LLM assistant using our verification-correct framework (yellow bar for the default mode and green bar for the uncertain mode). Accuracy (ranging from about 40.0 to 100.0) is plotted on the vertical axis, while each participator’s results are grouped along the horizontal axis. Comparing the Raw (pink) settings, H1’s baseline accuracy is clearly higher than that of H2 and H7. When the LLM assistant (yellow) or LLM assistant uncertain (green) is applied, the corrected results of H2 and H7 (they originally performed at a lower accuracy) show measurable improvements. Statistically significant p‐values shown in the same figure indicate that these gains are meaningful. Overall, our results illustrate that human participators with lower initial analytical performance benefited notably from LMM assistance, improving their answer accuracy by a large margin. This observation is further supported by comparing the performances between ARIEL and human assistants for correcting the answers from participators, shown in Figure 6 (c) and Extended Data Figure 8, which demonstrates that ARIEL has a generally better performance compared with the average level of human experts. Meanwhile, we show the improvement of using ARIEL as a corrector versus the raw answers via Sankey plots for H2 and H7. Based on Figure 6 (d), ARIEL can improve the accuracy of answers for over 65% questions with low initial accuracy, further supporting the strong capacity of ARIEL as a research assistant. Extended Data Figure 9 shows our consideration of inviting human evaluator using ARIEL to correct the previous answers, and the final accuracy is similar to the outputs from ARIEL, and thus we have demonstrated the potentials of using ARIEL as a research assistant for biology research.
@@
Figure 6 (e) shows a breakdown analysis over different categories of biomedical questions. Applying an LLM assistant over H2 or H7 consistently outperformed H2 and H7 for all categories, indicating the effectiveness of the LLM assistant for verifying and correcting H2 and H7 solutions. Applying an LLM assistant in the uncertainty mode where “I am not sure about the answer" is added after the human answer to deliberately encourage LLM to correct the answer outperforms the original H2 and H7 performance in most categories except for rank identification. Our LMM output inspection shows that LLM generates relevant context, but it struggles to consistently produce a final ranked answer for the rank identification task when the uncertainty mode is enabled. Among all categories, Insight Discovery exhibited the largest variance in accuracy across both human and LLM-assisted answers, consistent with earlier findings. This result highlights that generating novel biological insights remains a fundamental challenge, reflecting limitations not only in current LMMs but also in human expertise.
We present case studies in Figures 6 (f) and (g) to further illustrate the effectiveness of the verification–correction framework. In Figure 6 (f), the human participant failed to identify the target panel, and the LMM assistant also produced an incorrect response. However, unlike the human, the LMM provided explicit reasoning, demonstrating a more systematic attempt at problem solving. In Figure 6 (g), the human participant was unable to interpret the figure due to its complexity, whereas the LMM successfully retrieved relevant domain knowledge, identified and extracted the patterns in the “E" regions, and produced a correct explanation. Notably, human participants were permitted to consult external resources for this task, suggesting that the LMM exhibited a stronger ability to integrate genomic annotation knowledge in reaching the correct conclusion. These examples highlight two important points: (1) the LMM can correct human errors arising from knowledge gaps, and (2) the proposed verification–correction pipeline effectively elevates the performance of less analytically skilled participants to a level comparable with top human experts.
@@
Figure 7: Landscape and results of generating hypotheses from multimodal inputs. We report the average and standard error in all bar plots. (a) Our proposed framework of generating scientific hypotheses with two different modes, which all generate hypotheses. The HypoGenAgent mode has four different LMMs as group members, and their functions are explained in the main text. (b) Bar plot of evaluation results of different generated outputs with four metrics and their averaged score (Avg). (c) Examples of generated hypotheses with the same image inputs but from different modes. Important pathway and cellular activity information is boldfaced. The distinct information from different modes is underlined.
@@
Exploring the hypothesis-generation capacity of ARIEL framework with text-image paired inputs. Finally, we explore the potential of leveraging ARIEL to integrate multimodal understanding for scientific hypothesis generation, with the goal of inspiring novel directions in biomedical research. In this analysis, we reuse scientific images from the previous tasks but apply new prompts specifically designed for hypothesis generation. We investigate two modes: (1) a single-LLM mode using o1 as the base model, and (2) a Multi-LMM agent mode (HypoGenAgent), also based on o1. HypoGenAgent employs four distinct LMMs assigned to different roles, following a modified design from (29). Here the Analyst takes the multimodal inputs and perform quantitative analysis, the Engineer executes the experimental plan, the Scientist proposes new hypothesis, and the Criticist can verify the quality of generated hypothesis. An overview of the framework is provided in Figure 7 (a). To evaluate the quality of the generated hypothesis, we consider different aspects modified from (65) with a comprehensive cover for evaluating scientific insights:
@@
Each evaluation criterion was quantified on a scale from 0 to 100, where 0 indicates complete non-compliance and 100 indicates full compliance. To mitigate potential bias in the evaluation process, we employed Claude-Sonnet-3.7 as an automated judge (66), which scored the outputs from each generation mode (results shown in Figure 7 (b)). According to this figure, although HypoGenAgent (o1) has a higher score compared with o1 on average, we do not observe statistically significant differences in the hypothesis generated by two different modes (p-value>0.05>0.05 with two-sided Wilcoxon rank-sum test). To better understand the properties of the hypothesis generated by these modes, we select representative examples of generated hypotheses for in-depth analysis. Based on Figure 7 (c), we found that the hypothesis from o1 generally has a longer length and more figure-oriented information, while HypoGenAgent (o1) tends to generate hypotheses with shorter lengths and more general plans. Moreover, o1 tends to discuss more gene-function-level information while HypoGenAgent (o1)’s outputs contains therapeutic discovery. Therefore, there exists a completeness-conciseness balance when using ARIEL as a hypothesis generator, which may help researchers to design ideas that fit well with their requirements.
@@
  • With the help of verification and correction strategy, we can leverage the reasoning ability of advanced LMMs to help under-performing human participators by correcting their wrong answers, and thus improve their performances. Therefore, our experiment shows the potential of using LMMs as assistants for understanding complicated scientific figures.
@@ 3 Discussion
Furthermore, our results based on difficulty analysis imply that LLMs even have consistency with humans in their understanding of difficult problems. For example, LLMs also usually summarized abstract with low scores on the articles that most human participators marked as most difficult. Meanwhile, the answers for questions with the insight discovery category in the scientific figure understanding task also had larger variation compared with other categories, which was observed in both outputs from humans and outputs from models. These findings provide a new perspective on generation errors in foundation models. Finally, we demonstrated that ARIEL can also work as a hypothesis generator with multimodal inputs for accelerating scientific discovery, which aligns well with the perspectives of having agentic AI for accelerating scientific research (67).
@@ 4 Methods
Problem Statement. In this study, we focus on evaluating the outputs of LLMs understanding biomedical manuscripts and comparing the human- and LLMs-generated abstracts or answers. Here we define the participator $\mathcal{M}$ in the experiment as $O=\mathcal{M}(P,T,I)$, where $O$ represents outputs, $P$ represents prompts, $T$ represents text inputs, and $I$ represents image inputs. To evaluate the quality of LLMs-generated abstracts, $I$ is placed as empty, and we treat the original abstracts as ground truth $O^{T}_{G}$ and compare their differences by both quantitative approaches (denoted as evaluator $E(O_{G},O)$) and qualitative approaches. We also invite human experts to summarize the same testing documents and compare the differences from outputs. To evaluate the quality of LMMs-generated answers for figure understanding, $I$ is paired with question $P$, and $T$ is empty. Here we compare their outputs with human-annotated ground truth $O^{I}_{G}$ and invite human experts to answer the same questions and evaluate both human- and LMMs-generated outputs.
@@
Regarding the first part of the evaluation dataset, we download the open-access research manuscripts from PubMed (31) with a knowledge cut-off date later than 2024 January 1st and earlier than 2024 June 1st. The total dataset contains $n=2571$ papers, covering different topics. We further split the papers into training and testing datasets with a ratio of 0.7/0.3. The training dataset was used to fine-tune specific models, while the testing dataset was used to evaluate both open-source and closed-source models. For one paper, we took the abstract section out as ground truth, and deleted this section in the original paper as input texts. Our instruction was used to query the model to give us the summarization of targeted research manuscripts. Our evaluation does not suffer from the problems of data leakage, as our setting is harder and different from the generative pre-training stage. By removing the abstract from the context, tested LLMs cannot know directly infer the abstract based on the causality in text.
Regarding the second part of the evaluation dataset, we collected ten scientific figures from published manuscripts in top-tier journals focusing on genomics and proteomics research (23, 68, 69). We designed several questions to evaluate the performances of LMMs with different aspects, including figure summarization, color identification, model design, rank identification, recommendation, and insight discovery. We adjusted the prompt to provide necessary background information with questions as instruction and input texts, and combined the prompt with the paired figure as inputs of different LMMs. We generated the ground truth information of each question by referring to the original description of these figures in the manuscript. Our evaluation does not suffer from the problems of data leakage, as the questions are designed by ourselves without using AI models, and our selected figures have copyrights.
@@
For the long document summarization task, we consider seven different metrics for evaluation, including BLEU, ROUGE-1, ROUGE-2, ROUGE-L, BERT score (BERT), F1-Radgraph, and MEDCON, for evaluation (48, 49, 50, 51, 47). Details of the metrics are introduced below:
@@
5. MEDCON: MEDCON restricts the concepts and entities in QuickUMLS (70) semantic groups (Anatomy, Chemicals & Drugs, Device, Disorders, Genes & Molecular Sequences, Phenomena and Physiology) and we extract such sets from generated texts and ground truth texts. We then compute the F1 score based on these two sets. This score ranges from 0 to 1 and is later scaled to (0,100). A higher score means better model performance.
Based on the evaluation results recorded in (12), BERT and MEDCON scores have a high correlation with the correctness of outputs, while BLEU and ROUGE scores have a high correlation with the completeness of outputs. We compute the final score of each method by averaging the metric scores across all the testing manuscripts. The scores are scaled all into 0-100 and higher scores mean better model performances.
For the scientific figure understanding task, by following the design of (25, 23), we only consider the accuracy by comparing model outputs and ground truth information. Since we also need to consider the inference information as well as format information of model outputs, and thus our previous settings for long document summarization evaluation cannot be directly applied. We utilize Spacy (71) as a tool to split model outputs into different sentence blocks with similar methods and evaluate the correctness of each block. The accuracy is computed based on the number of correct blocks and the total number of blocks, for each question. For each figure, we compute the average score across different questions to report the performance of specific models in the given figure.
@@
LMM verification and correction to improve figure understanding capacity of humans. Scaling LMM test-time compute has been shown to be effective in improving the reasoning performance of LMMs on math and logical reasoning(72, 60, 73). We adopt the self-verification and self-correction component of SETS (72), a test-time scaling framework for iterative verification and correction and propose a verify-correct framework using an LMM to conduct verification and correction over human-written solutions for biological figure reasoning problems. In contrast with SETS which adopt self-verification and self-correction over LLM outputs, we conduct verification and correction over human outputs as well as iterative self-verification and self-correction over LMM-generated outputs.
@@
We use one of the most advanced general-purpose Multimodal reasoning models, OpenAI o1 (59) for evaluating LLM verification and correction over improving figure understanding capacity of humans. The temperature is set to 1 to promote the model to generate diverse and creative reasoning steps. The maximum iterations of verification and correction are set to 10 by following the general setting in this method (72) (In our task, observed minimal iteration is 6, and maximal iteration is 10). The workflow is illustrated in Algorithm 1.
@@
1:Input: System prompt $T_{S}$, question $Q_{M}$, scientific image $I_{S}$, human outputs $O_{H}$, verify prompt $T_{V}$, correct prompt $T_{C}$, number of iteration NN.
2:Helper Models: An LMM $\mathcal{M}(\cdot,\cdot,\cdot)$, concatenation function $\cdot||\cdot$.
@@
We propose two different modes for this process, including default mode and uncertain mode. The uncertain mode means we add “I am not sure about the answer." at the end of each human output to let the LMM agent be aware of the uncertainty level of such answers.